Three cadidates K, L and M were voted into office as school prefects. K secured 45% of the votes, L had 33% of the votes amd M had the rest of the votes. If M secured 1430, calculate,
a) the total number of votes cast
b) how many more votes K received thann L
THANK YOU

Answer by Cromlix(4381)

Hi there,
K secured 45% of the votes
L had 33% of the votes
M had the remainder which was:
100 - (45 + 33)
22%.
22% = 1430
1% = 1430/22
100% = (1430/22) x 100
100% = 6500 votes.
Total votes = 6500
K = 45% of 6500 = 2925 votes
L = 33% of 6500 = 2145 votes
K had (2925 - 2145) = 780 votes
more than L.
Hope this helps :-)